log9u = -3/2
Please solve with steps

Answer by stanbon(48569)

log9u = -3/2
Please solve with steps
----------------------
a = 9^(-3/2)
a = 3^(-3)
a = (1/3)^3
a = 1/27
===========
Cheers,
Stan H.